OpenStudy (secret-ninja):

Help. Medal. :) Check all polynomials that are perfect square trinomials: 9x^2 - 6x + 1 36b^2 - 24b + 8 16x^2 + 24x + 9 4a^2 - 10a + 25

Nnesha (nnesha):

hmmmmmmmm that ???????????????????

OpenStudy (alexandervonhumboldt2):

(a+b)^2=a^2+2ab+b^2 (a-b)^2=a^2-2ab+b^2

OpenStudy (alexandervonhumboldt2):

check if somebody fits that

OpenStudy (anonymous):

Try factoring them using this law:\[(x+y)(x+y)=x ^{2}+2xy+y\]\[(x-y)(x-y)=x ^{2}-2xy+y\]

OpenStudy (alexandervonhumboldt2):

A C

OpenStudy (secret-ninja):

Thanks you guys ^-^

OpenStudy (alexandervonhumboldt2):

k